OpenStudy (anonymous):

Simplify: (-3y)2(2x2) -6x2y2 12x3y 12x3y2 None of the above

OpenStudy (anonymous):

its very simple question!!!!!!!!!!

OpenStudy (mertsj):

\[(-3y)^2(2x^2)\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (skullpatrol):

Use the rule of exponents.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Have you used the Associative property before? It can be used both addition, multiplication, subtraction, but for multiplication, it states the following:\[a(b c)=(ab)c\] It just means that multiplication can be done in any order. Also, we need to know the following exponent law:\[(ab)^x=a^xb^x\] This just means that if we take values in a bracket to a power, then that can be re-written as all the numbers being taken to the same power. e.g. \((2*3)^2=2^2*3^2=4*9=36\). Applying rules we get:\[(-3y)^2(2x^2)=(-3^2y^2)(2x^2)=(9y^2)(2x^2)=(9*2)(x^2y^2)=18x^2y^2\] @InsanelyChaotic

OpenStudy (anonymous):

simplified?

OpenStudy (anonymous):

That's the final simplified form. It cannot be simplified further.

OpenStudy (anonymous):

-6x2y2 12x3y 12x3y2 None of the above

OpenStudy (anonymous):

Notice how I first used the exponent law on (-3y)^2 and the associative property to regroup the numbers so I can multiply 9 and 2 and x^2y^2 separately. And that gives me the answer.

OpenStudy (anonymous):

So what's the right answer then? @InsanelyChaotic

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Now may be try doing some questions yourself? That's the best way to learn. @InsanelyChaotic

OpenStudy (anonymous):

thank you, i did but i got stuck SO